OpenStudy (desmarie):

The molecular weight of CaCl2 is 110.98 g/mol. In order to make up 500.0 mL of a 0.500 M solution, you need 27.7 g. How is this true?

OpenStudy (desmarie):

@Photon336

OpenStudy (desmarie):

@radar

OpenStudy (photon336):

\[Molarity = \frac{ moles }{ L }\]

OpenStudy (photon336):

What we need to do is this we know what the molarity is and we know our volume right?

OpenStudy (desmarie):

well actually 27.7 is the answer but I dont know how they got that answer.

OpenStudy (photon336):

so we can easily check this by doing the following this is what we're going to do. \[molarity*volume*molar~mass = \frac{ moles }{ L }*L*\frac{ grams }{ mol }\]

OpenStudy (photon336):

\[0.5\frac{ moles }{ Liters }*0.5~Liters*110.98\frac{ grams }{ mole } = grams~CaCl_{2}\]

OpenStudy (desmarie):

ok I get how they go the 27.7 now thanks that helped a lot.

OpenStudy (photon336):

We've got to get to our answer too, take a look at what I did. essentially molarity is the number of moles per liter of solution. we want to find the number of grams of our compound in that solution. all we need to do is multiply the molarity by the volume in Liter to give us the moles of the compound. then the next step is easy.

OpenStudy (desmarie):

ok thanks I really appreciate it I wish I could give you more than 1 medal.

OpenStudy (photon336):

you can even try this out yourself. practice How many grams of NaCL are in a 0.25M solution when you know you have .2 Liters of of NaCl?

OpenStudy (desmarie):

would the answer be 2.3 g?

OpenStudy (photon336):

Show me your work

OpenStudy (desmarie):

0.2 moles/liters times 0.2 liters times 58.44 grams/mole = grams of NaCl